OpenStudy (anonymous):

graph the equation by plotting three points -x+2y=11

OpenStudy (anonymous):

can some please help me

OpenStudy (anonymous):

Solve for the format \[y = mx+b\] than substitute the x values 1, 2 and 3 into the linear equation. Solve for y and now you have a point (x, y) now you can plot it on a graph. 

OpenStudy (anonymous):

how do i change it to y=mx+b form

OpenStudy (anonymous):

first bring the x over to the right side.. -x +x +2y = x + 11 than you get 2y = x + 11 than divide both sides by 2 2y = x +11 -- ------ 2 2

OpenStudy (anonymous):

so 1y=2x+11/2

OpenStudy (anonymous):

Not quite... If you divide x by two what do you get ?

OpenStudy (anonymous):

0.5

OpenStudy (anonymous):

good 0.5x

OpenStudy (anonymous):

now you have a equation y = 0.5x + 11/2 can you put different numbers in for x and solve for y. ex y = 0.5(0) +11/2 y = 0 + 11/2 y = 11/2 *** Now I have a point (0 , 11/2) which I can plot on a graph.

OpenStudy (anonymous):

i cant plot 11/2 it had to be a whole number

OpenStudy (anonymous):

Than you can choose a point ex y = 0.5(1) + 11/2 y = 0.5+ 5.5 y = 6 *** than plot (1,6)

OpenStudy (anonymous):

ok can you please help me find x values and than i will do the y

OpenStudy (anonymous):

alright.. can you do x = 3 and x = -1

OpenStudy (anonymous):

3,7 and -1,5

OpenStudy (anonymous):

Good can you plot them ?

OpenStudy (anonymous):

yes thank you so much

OpenStudy (anonymous):

Nice rabaca.. well done =))

OpenStudy (anonymous):

thank you. for everything